OpenStudy (anonymous):

solve the trigonometric equation (sinx)(2cosx+1)=0 [0,360 degrees)

OpenStudy (anonymous):

i wouldn’t it is easier to solve for \(x\) when it is in factored form

OpenStudy (anonymous):

\[\sin(x)\left(2\cos(x)-1\right)=0\] \[\sin(x)=0\] or \[2\cos(x)+1=0\iff 2\cos(x)=-1\iff \cos(x)=-\frac{1}{2}\]

OpenStudy (campbell_st):

the easy method is to solve the factors sin(x) = 0 and 2cos(x) + 1 = 0 or cos(x) = -1/2 they will give the answers

OpenStudy (anonymous):

Im still not 100% sure what the answer is ?

OpenStudy (campbell_st):

ok... start with cos what angle gives cos(x) = 1/2  here is an exact value triangle